I need help solving this problem,
3/8(7+2x)+3 is greater than or equal to 18

Answer by RAY100(1637)

(3/8)(7+2x) +3 >= 18
.
(3/8) (7+2x) >= 15
.
Multiply by 8
.
3(7+2x) >= 120
.
21+6x >= 120
.
6x >= 99
.
x >= 16.5
.
Check,,,(3/8) (7+2(16.5)) +3 >= 18
.
(3/8)(40) +3 >= 18
.
15+3 >= 18,,,,,,ok